The Gauteng traffic department records show that 25% of all drivers wear seatbelts. In a random sample of 400 cars stopped at a roadblock in Gauteng, 152 of the drivers were wearing seatbelts. What is the probability that at most 18% of the drivers in Gauteng wear seatbelts? A 90% confidence interval for the proportion who wear seatbelts is given by
1
Expert's answer
2021-10-11T05:53:41-0400

Given "n=400, p=0.25, q=1-p=1-0.25=0.75."

"np=400(0.25)=100\geq10, nq=400(0.75)=300\geq10"

Then "\hat{p}" has approximately a normal distribution with "\mu_{\hat{p}}=p=0.25" and "\sigma_{\hat{p}}=\sqrt{\dfrac{pq}{n}}=\sqrt{\dfrac{0.25(0.75)}{400}}=\dfrac{\sqrt{3}}{80}."

Then


"P(\hat{p}<0.18)=P(Z<\dfrac{0.18-0.25}{\sqrt{3}/80})"

"\approx P(Z<-3.23316)\approx0.000612"

The sample proportion is computed as follows, based on the sample size "N = 400" and the number of favorable cases "X = 152:"


"\hat p = \displaystyle \frac{X}{N} = \displaystyle \frac{152}{400} = 0.38"

The critical value for "\alpha = 0.1" is "z_c = z_{1-\alpha/2} = 1.6449."

The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\\hat{p}-z_c\\sqrt{\\dfrac{\\hat{p}(1-\\hat{p})}{n}},\n\u200b\n\u200b"

"\hat{p}+z_c\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(0.38-1.6449\sqrt{\dfrac{0.38(1-0.38)}{400}},"

"0.38+1.6449\sqrt{\dfrac{0.38(1-0.38)}{400}})"

"=(0.34, 0.42)"

Therefore, based on the data provided, the "90\%"  confidence interval for the population proportion is "0.34 < p < 0.42," which indicates that we are "90\%" confident that the true population proportion "p" is contained by the interval "(0.34, 0.42)."